The length of a new rectangular playing field is 2 yards longer than double the width. If the perimeter of the rectangular playing field is 268 yards, what are its demensions?
Answer by ikleyn(52776)

.

    w + (2w + 2) = 268/2


    3w + 2 = 134.


You do the rest.


Completed.